# Masked Autoencoders are PDE Learners

### v2

@@ 4.4 PDE Feature Prediction
In 2D, the effects of fine-tuning is more pronounced. Within the pretraining set, only in the 2D Burgers task does freezing the MAE encoder outperform a supervised baseline; nevertheless, masked pretraining serves as a good initialization for supervised fine-tuning. In addition, despite differing physics, prior knowledge from simpler 2D PDEs seems to benefit regression on the Navier-Stokes equations. When classifying 2D Heat, Advection, and Burgers data based on their discretization, MAE models greatly benefit from pretraining on multi-resolution data. We hypothesize that in 2D PDEs, variable spatial resolutions can be challenging to distinguish due to flattening the spatial dimension when patchifying inputs, whereas in 1D PDEs the data is already flattened.
4.5 Conditional Time-stepping and Super-resolution
1D Experiments
To evaluate the use of the MAE encoder for practical tasks, we train a neural solver on various 1D PDEs to predict or upsample physics. In particular, we consider five PDEs (KdV-B, Heat, Burgers, Adv, KS) as well as two PDEs under varying boundary conditions (HeatBC, WaveBC) and predicting physics on various resolutions of the KdV-Burgers equation (Res.). For time-stepping, we incorporate temporal bundling and the pushforward trick to stabilize training [8]. Furthermore, we test on FNO [6] and Unet architectures [18, 62], and add conditioning information to hidden states after convolutions [63; 64]. For super-resolution (SR), we implement a pipeline in which a network encodes low-resolution physics before upsampling with a discretization inversion operator $D^{-1}$ (e.g. linear, bicubic interpolation) and mapping to an output functional space with a neural operator [50]. Following this, we implement a Resnet encoder [65; 66] followed by an interpolation scheme and FNO operator; both Resnet and FNO models are provided conditioning information from MAE encodings of low-resolution physics. For additional details, see Appendix D.
@@ 4.5 Conditional Time-stepping and Super-resolution
After pretraining an MAE encoder on the 1D KdV-Burgers equation, we compare the base neural solvers (FNO, Unet, SR) to conditioning on a frozen MAE embedding (-MAEf) and allowing the MAE encoder to fine-tune when conditioning (-MAE). Results in 1D are presented in Table 3. Within the pretraining distribution (KdV-B) and certain PDEs, MAE conditioning consistently improves time-stepping and super-resolution performance. In addition, allowing MAE encoders to fine-tune can further lower errors. However, there are various exceptions, in particular PDEs with unseen boundary conditions. Despite this, improvements are consistent across different neural solver architectures, suggesting that pretrained MAE models can be agnostic to downstream model choices.
@@
2D Experiments
Following the setup in 1D, we repeat time-stepping/super-resolution experiments on 2D PDEs (Heat, Adv, Burgers, NS) and a combined set of 2D Heat, Advection, and Burgers equations (Combined). Additionally, we evaluate time-stepping performance on the combined 2D Heat, Advection, and Burgers equations discretized at variable resolutions (Res.). We follow the same conditioning and training strategies as 1D experiments, but modify the architectures to support 2D inputs, and present results in Table 4. After pretraining an MAE encoder on the 2D Heat, Advection, and Burgers equations, we observe improvements in conditional physics prediction and upsampling. Improvements tend to be more pronounced in 2D; we hypothesize that the increased difficulty of the task increases the importance of MAE encoder guidance in time-stepping and super-resolution. This increased difficulty also manifests itself when extrapolating pretrained encoders to new PDEs, as the performance on the Navier-Stokes equations is limited. Nevertheless, we observe similar trends whereby MAE conditioning is agnostic to downstream architecture choices.
Conditioning Benchmarks
We consider two conditioning benchmarks: a randomly initialized ViT encoder that embeds PDE inputs to a conditioning vector as well as a linear model that encodes ground-truth coefficient or boundary condition information to a conditioning vector. We present detailed results in Appendix E.2 and E.3, and discuss the overall results here. In general, we observe that the randomly initialized, fine-tuned encoder also improves PDE prediction and upsampling, and this improvement generally matches or outperforms the performance of the frozen MAE encoder. However, allowing the MAE encoder to fine-tune generally outperforms this random initialization and approaches the linear benchmark.
@@
Limitations
Despite the work presented so far, we recognize limitations of the MAE encoder. It can be costly to fully fine-tune the pretrained encoder during time-stepping or super-resolution since it operates on all unmasked tokens, and as a result does not benefit from the speed gains during MAE pretraining. To address this, freezing the MAE encoder greatly improves the training speed but decreases performance, especially on unseen PDEs. With enough PDE data during pretraining, freezing MAE encoders can be more practical since the pretraining distribution would cover most downstream cases. Lastly, for simpler 1D dynamics when coefficient information is available, conditioning on ground-truth PDE parameters remains the best choice in most scenarios. However, these approaches are not exclusive; initial work suggests that models provided with both ground-truth information and MAE embeddings can outperform models provided with just one of either.
@@ 5 Conclusion
We have presented a new method that extends masked pretraining from vision and video domains to physics, and evaluated several PDE-specific modifications. We empirically validate MAE models through reconstructing a diverse set of 1D and 2D PDEs and show generalization behavior to different spatial discretizations and unseen equations. Furthermore, we evaluate the latent representations learned during MAE training and find structured trends despite training without labeled data. Furthermore, we propose operating in this latent space to interpolate between encoded PDE vectors. In practice, MAE encoders can also be used to improve PDE feature prediction, time-stepping, and super-resolution tasks across diverse physics scenarios. A promising direction would be to scale MAE models to larger datasets, as the current approach enjoys the same scalability as the originally proposed MAE [44]. Furthermore, future work could explore manipulating latent physics to generate new solutions or performing more sophisticated operations in this latent space.
Appendix A MAE Implementation and Ablation Studies
A.1 Implementation
To implement the MAE encoder and decoder, we use a ViT architecture [42], which uses a self-attention layer [46] and MLP, both with LayerNorms [67]. For MAE models evaluated in the paper trained on 1D and 2D PDEs, we present the pretraining parameters in Table 5(b). To study the effects of various hyperparameters, including model size, masking ratio, and patch size, we run ablation studies on masked reconstruction of 1D PDEs, and report reconstruction MSE errors on a validation set in Table 6(c). Overall, we find that increasing model size in limited data regimes—only 10000 KdV-Burgers samples were used in pretraining—tends to contribute to overfitting and increases validation errors. Predictably, increasing the masking ratio increases reconstruction errors as a result of less information being provided to MAE models. Furthermore, decreasing the patch size reduces errors but requires a higher computational cost, which is consistent with results in CV domains [54].
Following these design considerations, we use a MAE model with 5M params; pretraining with this setup takes around 3 hours on a NVIDIA GeForce RTX 2080 Ti GPU for the 1D KdV-Burgers dataset. In 2D, we use a MAE model with 10M params, which takes about 1 day to pretrain on a NVIDIA RTX A6000 GPU on the 2D Heat, Advection, and Burgers dataset.
@@ A.1 Implementation
A.2 Lie Point Symmetry Data Augmentations
We implement Lie Point Symmetry Data Augmentations according to Brandstetter et al. [2022b], including shifting and resampling PDE solutions with the Fourier shift theorem. Since we only augment PDE samples during pretraining, we consider symmetry groups for the 1D KdV-Burgers equation, as well as the 2D Heat, Advection, and Burgers equations. The 1D KdV-Burgers equation has the following symmetries, adapted from the generalized KdV-Burgers equation [68]:
$\partial_{t}u+\frac{j}{2}+\alpha u^{m}\partial_{x}u-\beta\partial_{xx}u+\gamma\partial_{xxx}u=0$ (1)
We consider the special case where $j=0,m=1$, resulting in the following Lie subalgebras [68]:
Taking the exponential map results in the following Lie groups [68]:
[…]
For the 2D Heat, Advection, and Burgers equations, there are many possible Lie subalgebras [68]. For simplicity, we only consider a basic subset of these that apply to all three equations, however, there is ample room to implement more symmetries:
These result in the following Lie groups in 2D:
[…]
Table 7: Lie augmentation ablation studies
Aug Ratio Error
0.00 2.48e-03
0.25 1.24e-03
0.50 1.17e-03
0.75 1.27e-03
1.00 1.37e-03
To measure the effect of augmentations on MAE reconstruction performance, we consider augmenting samples with different probabilities (augmentation ratio), with 0 corresponding to never augmenting samples and 1 corresponding to always augmenting samples. We report results on the 1D KdV-Burgers equation in Table 7. Lie augmentations consistently improve masked reconstruction on validation samples, even at low augmentation ratios. Additionally, the magnitude of the parameter $\epsilon$ can also modulate the strength of the augmentation; this is an additional consideration, but in practice we find that smaller values tend to work well.
Table 8: Multiresolution ablation studies
Strategy Error
None 4.40e-03
Pad 3.81e-03
Interp. 3.23e-03
Token 1.89e-03
A.3 Multi-resolution Details
To handle PDE inputs of varying discretizations, we consider the setup where inputs of shape ($n_{t},n_{x}$) or ($n_{t},n_{x},n_{y}$) can vary along their spatial dimension ($n_{x},n_{y}$) by multiples of the patch size ($p_{x},p_{y}$). During pretraining, snapshots in time, or time windows, are sampled with length $n_{t}$, which discretizes the PDE trajectories to a common time dimension. This is the same setup during fine-tuning; PDE features are predicted from an initial time window, and time-stepping/resolution experiments use temporal bundling [8].
Given a fixed patch size, varying spatial resolutions will affect the input sequence length to the ViT. This can be an issue for vanilla positional encodings, as the same position in the input sequence can now represent different spatial coordinates. This is especially pronounced in 2D; flattening two spatial dimensions results in positional information needing to wrap around inputs, and as a result, the same position can represent very different points in space when given different discretizations. To mitigate this issue, we propose strategies for embedding spatial resolution information into the ViT backbone.
The simplest strategy would be to pad the inputs to the maximum sequence length (Pad). However, this does not address the fact that positional information becomes overloaded and only provides information about input length. To address this, we consider interpolating positional embeddings from the maximum sequence length to variable sequence lengths (Interp.). This would allow tokens to receive approximate positions, however this would not be accurate for irregular grids. Lastly, to adapt to arbitrary grids, we consider using an embedder network, which could be a simple MLP or CNN, to tokenize spatial grids (Token). Validation MSE errors after training on varying resolutions of the 1D KdV-Burgers equation are reported in Table 8. We observe consistent improvement by using more sophisticated methods for embedding spatial discretizations.
In 1D, multiresolution pretraining is significantly easier to learn, so we use the interpolation strategy for simplicity. In 2D, the use of tokenized spatial grid information becomes more important, so we use this strategy for 2D multiresolution experiments. The hyperparameters and training costs are the same as those presented in Appendix A.1.
Appendix B PDE Details
In this section we detail the various PDEs considered in the paper. For convenience, we copy the equations from Section 4.1 here as well.
B.1 1D PDEs
1D KdV-Burgers equation
The KdV-Burgers equation is a diverse equation that contains diffusive, shock formation, and dispersive behaviors:
$\partial_{t}u+\alpha u\partial_{x}u-\beta\partial_{xx}u+\gamma\partial_{xxx}u=\delta(t,x)$ (10)
We use a varied form that contains the Heat, Burgers, and KdV equations as corner cases modulated by the coefficient values ($\alpha,\beta,\gamma$), and is solved numerically according to the setup in Brandstetter et al. [2022a] with periodic boundary conditions. Both initial conditions and forcing terms are generated from the periodic function $\delta(x,t)$, where we uniformly sample $A_{j}\in[-0.5,0.5],\omega_{j}\in[-0.4,0.4],l_{j}\in\{1,2,3\},\phi_{j}\in[0,2\pi)$ while fixing $J=5,L=16$.
To generate diverse PDE samples, the coefficients are sampled uniformly from $\alpha\in[0,3],\beta\in[0,0.4],\gamma\in[0,1]$. Furthermore, samples are generated with a discretization $(n_{t},n_{x})=(250,100)$ on an interval $x=[0,16]$ from $t=0$ to $t=2$.
1D Heat and Burgers equations
The 1D Heat and Burgers equations describe the common phenomena of diffusion and shock formation:
$\partial_{t}u-\nu\partial_{xx}u=\delta(t,x),\quad\text{Heat}$ (12)
$\partial_{t}u+u\partial_{x}u=\delta(t,x),\quad\text{Burgers}$ (13)
We solve the equations with the same periodic BCs, initial conditions, and forcing function setup and as the 1D KdV-Burgers equation [8], but by setting the appropriate coefficient values. Specifically, from Equation 10 we uniformly sample $\nu=\beta\in[0.1,0.8]$ for the Heat equation and fix $\alpha=0.5,\beta=0$ to model the inviscid Burgers equation. These equations are also solved with a discretization $(n_{t},n_{x})=(250,100)$ on an interval $x=[0,16]$ from $t=0$ to $t=2$.
To generate data for the Heat equation that enforces Dirichlet or Neumann boundary conditions, we write the Heat equation in its variational form after an implicit Euler discretization:
[…]
This formulation can be solved using FEniCS [69, 70]. To simplify the boundary value problem, we set the forcing term $f=0$. Furthermore, we set [$u(x=0)=u(x=L)=0$] to enforce Dirichlet BCs, and [$\partial_{x}u(x=0)=\partial_{x}u(x=L)=0$] to enforce Neumann BCs; however, in both of these cases, the initial conditions in Equation 11 need to be modified to respect the new BCs.
1D Advection Equation
The 1D Advection equation is a simple model of transport phenomena:
$\partial_{t}u+c\partial_{x}u=0$ (17)
We solve the Advection equation with periodic BCs using its analytical solution:
$u(x,t)=u_{0}(x-ct)$ (18)
Initial conditions are generated according to Equation 11 and the wave speed is uniformly sampled from $c\in[0.1,5]$. The solution domain and discretization are the same as previous cases, with $(n_{t},n_{x})=(250,100)$, $x=[0,16]$, and time ranging from $t=0$ to $t=2$.
1D Wave Equation
The 1D Wave equation is a common equation that describes wave propagation across many disciplines:
$\partial_{tt}u-c^{2}\partial_{xx}u=0$ (19)
We solve the equation according to the setup in Brandstetter et al. [2022a] with Dirichlet ($u(x=0)=u(x=L)=0$) and Neumann ($\partial_{x}u(x=0)=\partial_{x}u(x=L)=0$) BCs, resulting in waves that either bounce or reflect off boundaries. The wave speed is fixed at $c=2$, and the initial condition is a Gaussian pulse with unit amplitude and with its peak randomly sampled on the spatial domain. Lastly, the equation is solved from $t=0$ to $t=100$ on the interval $x=[-8,8]$ with a discretization $(n_{t},n_{x})=(250,100)$.
1D Kuramoto-Sivashinsky Equation
The 1D KS equation is a chaotic fourth-order PDE with rich, complex dynamics:
$\partial_{t}u+u\partial_{x}u+\nu\partial_{xx}u+\partial_{xxxx}u=0$ (20)
We numerically solve this equation with periodic BCs according to Brandstetter et al. [2022b], however, the initial conditions are modified from the original work to match Equation 11. Following the data setup proposed by Lippe et al. [2023], we additionally modify the original work to uniformly sample $\nu\in[0.75,1.25]$ to vary the second-order term in the KS equation. Furthermore, due to the unique dynamics of the KS equation, we solve the PDE from $t=0$ to $t=100$ on the interval $x=[0,64]$ with a discretization of $(n_{t},n_{x})=(100,100)$.
B.2 2D PDEs
2D Heat, Advection, and Burgers Equations
The 2D Heat, Advection, and Burgers are simple models of diffusion, transport, and shock phenomena in 2D:
$\partial_{t}u-\nu\nabla^{2}u=0,\quad\text{Heat}$ (21)
$\partial_{t}u+\mathbf{c}\cdot\nabla u=0,\quad\text{Advection}$ (22)
$\partial_{t}u+u(\mathbf{c}\cdot\nabla u)-\nu\nabla^{2}u=0\quad\text{Burgers}$ (23)
We use a 1st-order central differences scheme to discretize the diffusive term $(\nabla^{2}u)$, and a 1st-order upwinding scheme to discretize the nonlinear convective term $(u\nabla u)$. Additionally, the 2D Advection equation is solved with its analytical solution:
$u(x,y,t)=u_{0}(x-c_{x}t,y-c_{y}t)$ (24)
[…]
2D Incompressible Navier-Stokes Equations
The 2D incompressible Navier-Stokes equations are an important set of PDEs that describe viscous fluid flow. Following the setup in Li et al. [2020], we formulate the NS equations in vorticity form with variable viscosity and forcing function amplitude:
$\partial_{t}\omega+\mathbf{u}\cdot\nabla\omega-\nu\nabla^{2}\omega=f(x,y),\quad\nabla\cdot\mathbf{u}=0,\quad\nabla\times\mathbf{u}=\omega$ (26)
$f(x,y)=A(sin(2\pi(x+y))+cos(2\pi(x+y)))$ (27)
The solution is solved on a grid $(n_{t},n_{x},n_{y})=(100,64,64)$ on a solution domain $(x,y)=[0,1]^{2}$ from $t=0$ to $t=25$. PDE parameters are uniformly sampled from $\nu\in\{\{1,2,3,4,5,6,7,8,9\}\times 10^{-\{6,7,8,9\}}\}$ and $A\in\{1,2,3,4,5,6,7,8,9,10\}\times 10^{-3}$. Lastly, initial conditions $\omega_{0}$ are sampled according to Li et al. [2020] from a Gaussian random field.
Appendix C Additional MAE Examples
C.1 Additional 1D MAE Predictions

### v3

@@ 5.4 Conditional Time-stepping and Super-resolution
For super-resolution (SR), we implement a pipeline in which a network encodes low-resolution physics before upsampling with a discretization inversion operator $D^{-1}$ (linear interpolation in 1D and bicubic interpolation in 2D) and mapping to an output function space with a neural operator (Yang et al., 2023b).Following this, we implement a Resnet encoder (Wang et al., 2021b; Zhang et al., 2018) followed by an interpolation scheme and FNO operator; both Resnet and FNO models are provided conditioning information from MAE encodings of low-resolution physics. The motivation behind using a two-step SR pipeline is to learn a vector embedding using the Resnet, then map from vector to function space with $T^{-1}$, and finally transform this function using an learnable operator to the output. For additional details on hyperparameters, see Appendix C. Additionally, we evaluate a simple baseline using linear interpolation to upsample low-resolution inputs (Interp.).
@@
After pretraining an MAE encoder on the 1D KdV-Burgers equation, we compare the base neural solvers (FNO, Unet, SR) to conditioning on a frozen MAE embedding (-MAEf) and allowing the MAE encoder to fine-tune when conditioning (-MAE). Results in 1D are presented in Table 5. Within the pretraining distribution (KdV-B) and certain PDEs, MAE conditioning consistently improves time-stepping and super-resolution performance. In addition, allowing MAE encoders to fine-tune can further lower errors. However, there are various exceptions, in particular PDEs with unseen boundary conditions. Despite this, improvements are consistent across different neural solver architectures, suggesting that pretrained MAE models can be agnostic to downstream model choices. In addition, in 1D, SR results are less significant suggesting that simple interpolation schemes are often enough for these phenomena, especially for simple equations such as the advection or wave PDEs.
2D Experiments
Following the setup in 1D, we repeat time-stepping/super-resolution experiments on 2D PDEs (Heat, Adv, Burgers, NS) and a combined set of 2D Heat, Advection, and Burgers equations (Combined). Additionally, we evaluate time-stepping performance on the combined 2D Heat, Advection, and Burgers equations discretized at variable resolutions (Res.). We follow the same conditioning and training strategies as 1D experiments, but modify the architectures to support 2D inputs, and present results in Table 6. Additionally, the interpolation baseline implements bicubic upsampling (Interp.). After pretraining an MAE encoder on the 2D Heat, Advection, and Burgers equations, we observe improvements in conditional physics prediction and upsampling. Improvements tend to be more pronounced in 2D; we hypothesize that the increased difficulty of the task increases the importance of MAE encoder guidance in time-stepping and super-resolution. However, out-of-distribution datasets are still challenging: when extrapolating pretrained encoders to new PDEs, such as the Navier-Stokes equations, the performance is limited. Nevertheless, we observe similar trends whereby MAE conditioning is agnostic to downstream architectures.
Additional Benchmarks
We consider two additional benchmarks: a randomly initialized ViT encoder that embeds PDE inputs to a conditioning vector as well as a linear model that encodes ground-truth coefficient or boundary condition information to a conditioning vector. We present detailed results in Appendix E.2 and E.4, and discuss the overall results here. In general, we observe that the randomly initialized, fine-tuned encoder also improves PDE prediction and upsampling, and this improvement generally matches or outperforms the performance of the frozen MAE encoder. However, allowing the MAE encoder to fine-tune generally outperforms this random initialization and approaches the linear benchmark.
@@
Lastly, to benchmark our method against other pretraining methods, we consider pretraining an encoder using a contrastive self-supervised technique proposed by Mialon et al. (2023), which relies on using Lie symmetries to cluster PDE data in a learned latent space. We contrastively pretrain an encoder on 1D KdV-Burgers data and the evaluate conditional timestepping performance on various downstream 1D PDEs. We present these results in Appendix E.3. To summarize, our approach is on par with Lie contrastive pretraining for PDE samples within the pretraining distribution; however, when extrapolating to unseen PDEs, masked pretraining is able to outperform contrastive methods.
6 Discussion
In this section, we discuss results and provide additional insight. Although masking as a pretraining strategy is not physically valid, this requirement does not seem to be necessary for learning. Both in this study, and in related works (Hao et al., 2024; Zhou et al., 2024; Rahman et al., 2024), noising or masking strategies are used to improve model performance, which both represent artificial phenomena. Within the context of masking, it is also natural to ask if a unique solution exists given a masked input. Certainly at the extreme where all of the input is masked the solution is not unique; however, it seems empirically that only a small amount of information (25% in 1D and 10% in 2D) is needed, which is corroborated by related work in the CV domain (Feichtenhofer et al., ). We can observe this through the validation error: if this quantity is small, a unique solution is being regressed in the validation set.
While promising, the presented results in time-stepping and super-resolution would likely not outperform a foundation model trained specifically for these tasks (Hao et al., 2024; Herde et al., 2024). This is because direct transfer learning has been shown to outperform surrogate objectives when pretraining for PDEs (Zhou et al., 2024); one hypothesized reason for this is the lack of abundant unlabeled data in the PDE domain (or equivalently, the downstream task uses also unlabeled data). However, encoder-style approaches such as this work or contrastive PDE encoders (Mialon et al., 2023; Zhang et al., 2023b) are more flexible than foundation models, capable of being applied to arbitrary downstream architectures and different fine-tuning tasks. Additionally, when using a surrogate objective to train an encoder, models can learn more general latent representations, compared to using a neural solver or foundation model to only predict the next step of a PDE rollout. We show some preliminary results demonstrating this in Appendix F. While this versatility is interesting, the practicality of this is unclear, since time-stepping is so singularly important. However, this research direction is still underexplored and perhaps future work will find an interesting set of uses for PDE encoders.
Within the context of masked pretraining, there are a few additional limitations to be recognized. It can be costly to fully fine-tune the pretrained encoder during time-stepping or super-resolution since it operates on all unmasked tokens, and as a result does not benefit from the speed gains during MAE pretraining. This makes full fine-tuning likely infeasible when compared to baselines. To address this, freezing the MAE encoder greatly improves the training speed but decreases performance, especially on unseen PDEs. With enough PDE data during pretraining, freezing MAE encoders can be more practical since the pretraining distribution would cover most downstream cases. Lastly, for simpler 1D dynamics when coefficient information is available, conditioning on ground-truth PDE parameters remains the best choice in most scenarios. However, these approaches are not exclusive; initial work suggests that models provided with both ground-truth information and MAE embeddings can outperform models provided with just one of either.
@@ 7 Conclusion
We have presented a new method that extends masked pretraining from vision and video domains to physics, and evaluated several PDE-specific modifications. We empirically validate MAE models through reconstructing a diverse set of 1D and 2D PDEs and show limited generalization behavior to different spatial discretizations and unseen equations. Furthermore, we evaluate the latent representations learned during MAE training and find structured trends that can be used to predict PDE features. In practice, MAE encoders can also be used to improve time-stepping and super-resolution tasks across diverse physics scenarios.
A promising direction would be to scale MAE models to larger datasets, as the current approach exhibits the same scalability as the originally proposed MAE (He et al., 2021). Additionally, future work could explore masked modeling approaches in more complex 2D and 3D problems. Lastly, future work could explore manipulating latent physics to generate new solutions or performing arithmetic in an autoencoder latent space. We present a potential setup for this in Appendix D which relies on encoding solutions from separate equations, adding them in latent space, and decoding this latent vector.
Appendix A MAE Implementation
To implement the MAE encoder and decoder, we use a ViT architecture [Dosovitskiy et al., 2020], which uses a self-attention layer [Vaswani et al., 2023] and MLP, both with LayerNorms [Ba et al., 2016]. We present hyperparameters in Table 7(b). To study the effects of various hyperparameters, including model size, masking ratio, and patch size, we run ablation studies on masked reconstruction of 1D PDEs, and report reconstruction MSE errors on a validation set in Table 8(c). Overall, we find that increasing model size in limited data regimes—only 10000 KdV-Burgers samples were used in pretraining—tends to contribute to overfitting and increases validation errors. Predictably, increasing the masking ratio increases reconstruction errors as a result of less information being provided to MAE models. Furthermore, decreasing the patch size reduces errors but requires a higher computational cost, which is consistent with results in CV domains [Beyer et al., 2023].
In 1D, the MAE is trained on a single NVIDIA GeForce RTX 2080 Ti, and reaches convergence in about 6 hours. In 2D, the MAE is trained on a single NVIDIA RTX A6000, and reaches convergence in about 24 hours. Model size, masking ratio, and patch size all affect the computational cost and can be used to tradeoff performance for compute and memory.
To motivate the ViT architecture, we investigate the effect of different model choices and hyperparameters on performance. We find that FNO autoencoders tend to have poor reconstruction capabilities due to introducing spurious high-frequency modes when masking spatially. Unet approaches fare better but still suffer sharp boundaries across masked and unmasked regions. Additionally, we evaluate different ViT variants, such as ViViT (axial attention) and Swin Transformer (window attention). We find that restricting the attention mechanism reduces performance, and the additional speedups were not significant since masking out large portions of the input already reduces computation. Lastly, model performance tends to vary smoothly with changes in hyperparameters; for example, reducing patch size slightly increases performance across downstream tasks.
@@ Appendix A MAE Implementation
Appendix B Additional MAE Examples
B.1 Additional 1D MAE Predictions